OpenStudy (anonymous):

Could someone help me with this one..? x y 2 0 4 4 6 8 Find the average rate of change for the given function from x = 2 to x = 4. −2 negative one-half one half 2

OpenStudy (anonymous):

x= 2, 4, 6 y=0, 4, 8

OpenStudy (isaiah.feynman):

Please next time, write this in mathematical form. I can't decipher your question.

OpenStudy (anonymous):

oops. sorry :///

OpenStudy (anonymous):

its a table and the x is 2, 4, and 6 and the y is 0, 4, 8

OpenStudy (anonymous):

and it's asking what the rate of change is if it's -2, -1/2, -1/2, and 2

jimthompson5910 (jim_thompson5910):

when x = 2, what is y

OpenStudy (anonymous):

0

jimthompson5910 (jim_thompson5910):

when x = 4, what is y

OpenStudy (anonymous):

lol right..? :D

OpenStudy (anonymous):

4!

OpenStudy (anonymous):

x = 6 and y = 8

jimthompson5910 (jim_thompson5910):

good, we don't need to worry about the x = 6 and y = 8 pair

jimthompson5910 (jim_thompson5910):

the difference from x = 2 to x = 4 is 2 units (4 - 2 = 2) the difference from y = 0 to y = 4 is 4 units (4 - 0 = 4) divide the two differences to get (y difference)/(x difference) = 4/2 = 2

jimthompson5910 (jim_thompson5910):

another way to think of it: the average rate of change is the slope of the line through the two points (2,0) and (4,4)

OpenStudy (anonymous):

so my answers 2!

jimthompson5910 (jim_thompson5910):

yes it is

OpenStudy (anonymous):

THANK YOU SOOO Much!! you're amazing :D

jimthompson5910 (jim_thompson5910):

you're welcome